# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v4

@@ Appendix D A Monte Carlo study
This appendix present a Monte Carlo study that assesses the performance of the proposed penalized maximum likelihood estimator and the three-step estimation procedure discussed in Section 4. We consider simple bivariate (structural) TVAR and LSTVAR models with $p=1$ and $M=2$, two specifications of each, with the first lag of the first variable as the switching variable. To maintain some constancy, we use the same impact matrices across all specifications, $\text{vec}(B_{1})=(0.6,-0.3,0.2,0.4)$ and $\text{vec}(B_{2})=(0.7,0.1,0.3,0.8)$, as well as the same degrees-of-freedom and skewness parameter values, $\nu_{1}=2.5$, $\nu_{2}=12.0$, $\lambda_{1}=-0.5$, and $\lambda_{2}=0.2$. For TVAR models, we use the threshold parameter value $r_{1}=0.8$ and for LSTVAR models the location and scale parameter values $c=0.8$ and $\gamma=5.0$.
@@
For both the TVAR model and the LSTVAR model, we consider two specifications of the AR matrices, presented in Table D.1. In the first specification (labelled as TVAR 1 and LSTVAR 1 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the stability region (see Section 4.1). In the second specification (labelled as TVAR 2 and LSTVAR 2 for the TVAR and LSTVAR models, respectively), the AR matrices are well inside the region where penalization term is strictly positive, allowing us to assess the performance of our estimation method when the true parameter value lies in the penalization region. To illustrate the closeness to the boundary, the modulus of the roots of the companion form AR matrix are presented in Table D.1 (as in our empirical application, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$ for the penalization term, so the penalization term is strictly positive when at least one the roots of the companion form AR matrix is larger than $0.95$ in modulus). The intercept parameter values are adjusted so that the unconditional means of the regimes are approximately equal in both of the specifications (see Table D.1), so that the main difference between the specifications is that the second one includes true parameter values from inside the penalization region.
@@
$\phi_{1}$ $A_{1,1}$ $\mu_{1}$ $||\boldsymbol{A}_{1}||$ $\phi_{2}$ $A_{2,1}$ $\mu_{2}$ $||\boldsymbol{A}_{2}||$
[…]
@@
Table D.1: The parameters values of $\phi_{1},\phi_{2},A_{1,1}$, and $A_{2,1}$ used for the models TVAR 1, TVAR 2, LSTVAR 2, and LSTVAR 2 in the Monte Carlo study. Also, the corresponding unconditional means of the regimes $\mu_{m}=(I_{d}-A_{m,1})^{-1}\phi_{m}$, $m=1,2$, and modulus of the eigenvalues of the companion form AR matrices of the regimes $||\boldsymbol{A}_{m}||$, $m=1,2$.
@@
For each of the four model specifications, we generate $500$ samples of length $250,500,1000,2000$, and $10000$. Then, we fit the first-order two-regime TVAR model (for specifications TVAR 1 and 2) or LSTVAR model (for specifications LSTVAR 1 and 2) to each of the generated sample using the three-step estimation procedure based on the penalized log-likelihood function, proposed in Section 4. After estimation, the ordering and signs of the columns of $B_{1}$ and $B_{2}$ are standardized by assuming, without loss of generality, that the degrees-of-freedom parameters are in an ascending order and that the first skewness parameter is positive and the second one negative, in line with the true parameter values. To estimate the bias of the PML estimator, we calculate the average estimation errors $\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}$ over the $500$ Monte Carlo repetitions, and to estimate the spread of the estimates, we calculate their sample standard deviations.
$T=250$ $T=500$ $T=1000$ $T=2000$ $T=10000$
TVAR 1 $\phi_{1}$ $0.01$ $(0.07)$ $0.01$ $(0.05)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$0.00$ $(0.07)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$\phi_{2}$ $0.00$ $(0.16)$ $-0.01$ $(0.10)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.02)$
$0.01$ $(0.28)$ $0.00$ $(0.19)$ $0.00$ $(0.13)$ $0.00$ $(0.09)$ $0.00$ $(0.04)$
$A_{1,1}$ $-0.01$ $(0.08)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$-0.01$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
$-0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$A_{2,1}$ $-0.01$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$-0.01$ $(0.11)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$
$0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
$-0.01$ $(0.07)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$B_{1}$ $0.20$ $(0.52)$ $0.09$ $(0.33)$ $0.05$ $(0.21)$ $0.02$ $(0.12)$ $0.01$ $(0.04)$
$-0.08$ $(0.34)$ $-0.04$ $(0.19)$ $-0.02$ $(0.11)$ $-0.01$ $(0.06)$ $0.00$ $(0.02)$
$-0.02$ $(0.09)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$-0.03$ $(0.10)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$B_{2}$ $0.30$ $(0.61)$ $0.14$ $(0.39)$ $0.07$ $(0.24)$ $0.03$ $(0.13)$ $0.00$ $(0.04)$
$0.04$ $(0.15)$ $0.02$ $(0.08)$ $0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$-0.01$ $(0.07)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
$-0.03$ $(0.18)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$r_{1}$ $0.00$ $(0.07)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$ $0.00$ $(0.00)$
$\nu_{1}$ $0.05$ $(0.53)$ $0.02$ $(0.34)$ $0.01$ $(0.23)$ $0.00$ $(0.16)$ $0.00$ $(0.07)$
$\nu_{2}$ $165.03$ $(700.13)$ $105.25$ $(757.09)$ $13.98$ $(174.19)$ $1.48$ $(6.32)$ $0.16$ $(1.40)$
$\lambda_{1}$ $-0.03$ $(0.09)$ $-0.01$ $(0.06)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$\lambda_{2}$ $0.01$ $(0.10)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
TVAR 2 $\phi_{1}$ $-0.02$ $(0.12)$ $-0.01$ $(0.09)$ $-0.01$ $(0.06)$ $-0.01$ $(0.05)$ $0.00$ $(0.02)$
$-0.05$ $(0.14)$ $-0.04$ $(0.11)$ $-0.02$ $(0.07)$ $-0.02$ $(0.04)$ $-0.01$ $(0.02)$
$\phi_{2}$ $-0.02$ $(0.16)$ $-0.01$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.02)$
$-0.01$ $(0.30)$ $0.00$ $(0.18)$ $0.01$ $(0.12)$ $0.01$ $(0.08)$ $0.00$ $(0.04)$
$A_{1,1}$ $-0.02$ $(0.07)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.00)$
$-0.01$ $(0.07)$ $-0.01$ $(0.05)$ $-0.01$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$0.00$ $(0.07)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$-0.03$ $(0.06)$ $-0.02$ $(0.04)$ $-0.01$ $(0.03)$ $-0.01$ $(0.02)$ $-0.01$ $(0.01)$
$A_{2,1}$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
$0.00$ $(0.10)$ $0.00$ $(0.06)$ $0.00$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$-0.01$ $(0.04)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$
$-0.02$ $(0.07)$ $-0.01$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$B_{1}$ $0.38$ $(0.64)$ $0.16$ $(0.40)$ $0.08$ $(0.24)$ $0.04$ $(0.12)$ $0.01$ $(0.04)$
$-0.15$ $(0.43)$ $-0.07$ $(0.25)$ $-0.04$ $(0.15)$ $-0.02$ $(0.08)$ $0.00$ $(0.03)$
$-0.04$ $(0.12)$ $-0.02$ $(0.08)$ $-0.01$ $(0.04)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
$-0.05$ $(0.18)$ $-0.02$ $(0.07)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
$B_{2}$ $0.55$ $(0.73)$ $0.24$ $(0.44)$ $0.12$ $(0.25)$ $0.05$ $(0.12)$ $0.01$ $(0.04)$
$0.06$ $(0.17)$ $0.03$ $(0.09)$ $0.01$ $(0.05)$ $0.01$ $(0.03)$ $0.00$ $(0.01)$
$-0.01$ $(0.13)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
$-0.05$ $(0.25)$ $-0.01$ $(0.08)$ $0.00$ $(0.02)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$
$r_{1}$ $0.08$ $(0.45)$ $0.02$ $(0.24)$ $0.00$ $(0.02)$ $0.00$ $(0.01)$ $0.00$ $(0.00)$
$\nu_{1}$ $0.05$ $(2.77)$ $-0.07$ $(0.29)$ $-0.05$ $(0.21)$ $-0.03$ $(0.15)$ $0.00$ $(0.07)$
$\nu_{2}$ $107.17$ $(542.02)$ $93.32$ $(1273.27)$ $10.09$ $(81.02)$ $1.29$ $(5.10)$ $0.09$ $(1.47)$
$\lambda_{1}$ $-0.04$ $(0.11)$ $-0.02$ $(0.05)$ $-0.01$ $(0.04)$ $-0.01$ $(0.03)$ $0.00$ $(0.01)$
$\lambda_{2}$ $-0.01$ $(0.12)$ $0.00$ $(0.07)$ $0.00$ $(0.05)$ $0.00$ $(0.03)$ $0.00$ $(0.01)$
Table D.2: Results for the structural TVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-step estimation method based on the PML estimator discussed in Section 4. The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses. The results are based on $500$ Monte Carlo repetitions.
@@
Table D.3: Results for the structural LSTVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-step estimation method based on the PML estimator discussed in Section 4. The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses. The results are based on $500$ Monte Carlo repetitions.
@@
The results of the Monte Carlo study are presented in Table D.2 for the TVAR models and in Table D.3 for the LSTVAR models. The results show that there is some but mostly very small bias, which diminishes when the sample size increases. Since also the standard deviation of the estimates diminishes with the sample size, our results are in line with consistency of the PML estimator. Nonetheless, for certain parameters, there appears to be a notable bias in small samples.
@@
First, the estimator of the second degrees-of-freedom parameter (true parameter value $\nu_{2}=12$) is substantially biased upwards in samples of length $T=250,500$, and $1000$, and also the standard deviation of the estimates is very large. This can explained by the property of the skewed $t$-distribution that when the degrees-of-freedom parameter value is already large, increasing it more has virtually no effect on the shape of the distribution. Consequently, if a large degrees-of-freedom parameter fits the data, the numerical optimization algorithms are incapable of discriminating between large and very large parameter values. Hence, possibly very large degrees-of-freedom parameter values drawn by random in Step 2 estimation may remain large through Step 3 estimation. Since it is clear that there is virtually no difference in the economic interpretations of large (say, $100$) and very large (say, $1000$) degrees-of-freedom parameter values, the large upward bias should not be a concern.
@@
Second, in the LSTVAR models, there is a substantial upward bias also for the scale parameter $\gamma$ when the sample size is small. A large scale parameter value indicates almost discrete regime-switches, suggesting that the PML estimator is, in small samples, biased towards estimating too fast regime-switches when they are smooth in the true data generating process ($\gamma=5$). The upward bias is possibly heightened by the property that when $\gamma$ is large, the regime-switches are already nearly discrete, making it difficult for numerical estimation algorithms to properly distinguish between large and very large values, particularly in small samples.
@@
Third, the estimators of the first elements of the first rows of $B_{1}$ and of $B_{2}$ exhibit a notable upward bias when the sample is small. A possible explanation is that the bias is related to the heavy tails of the skewed $t$-distribution, as the true degrees-of-freedom parameter value is small ($\nu_{1}=2.5$), and also, the estimators of the first elements of the second rows of $B_{1}$ and $B_{2}$ are biased more than the estimators of most of the other parameters. In particular, the estimates may be sensitive to extreme values drawn for the heavy-tailed skewed $t$ shocks, potentially amplifying small-sample bias.
@@
In general, the small-sample bias of the PML estimator seems to be slightly larger the specifications where the true AR matrices are in the penalization region (TVAR 2 and LSTVAR 2) than the specifications where the true AR matrices lie well inside the stability region (TVAR 1 and LSTVAR 1). The differences are, however, only small, and even in small samples, there does not appear to be any particular issues with the estimation accuracy of the AR matrices lying inside the penalization region. Also, the TVAR models seem to be slightly more accurately estimated than the LSTVAR models, possibly because, unlike in TVAR models, in LSTVAR models also the smoothness of regime-switches needs to be estimated.
Appendix E Details on the empirical application
@@
The estimation of the models, residual diagnostics, and computation of the GIRFs are carried out with the accompanying R package sstvars (Virolainen, 2025a) in which the introduced methods have been implemented. The R package sstvars also contains the dataset studied in the empirical application. In the below residual based model diagnostics, we have not labelled the residual series by any of the variables, because labelling them effectively requires labelling the corresponding structural shocks, and only the CPU shock has been labelled.
@@
To study whether our model adequately captures the autocorrelation structure of the data, we depict the sample autocorrelation functions (ACF) and cross-correlation functions (CCF) of the (non-standardized) residuals for the first $24$ lags in Figure E.1. There is not much autocorrelation left in the residuals, but there is a particularly large correlation coefficient (CC) in the lag zero CCFs between the fourth and fifth residual series. In addition, there are a few moderately sized CCs in the ACFs and CCFs of some of the residual series, but the vast majority of the CCs are small. Hence, with the exception of the single large CC in the lag zero CCFs of the fourth and fifth residual series, our LSTVAR model appears to capture the autocorrelation structure of the data reasonably well.
@@
To study our model’s adequacy to capture the conditional heteroscedasticity in the data, we depict the sample ACFs and CCFs of squared standardized residuals for the first $24$ lags in Figure E.2. There is a very large CC in the lag zero CCFs between the second and third residual series, which, however, disappears if the COVID-19 residual from March 2020 is excluded (not shown). In addition, there are several sizeable CCs in the ACFs and CCFs of some of the residual series. Thus, there are some inadequacies in capturing the conditional heteroscedasticity, but with the exception of the large CC in CCFs between the second and third residual series (due to the extreme COVID-19 volatility), the inadequacies do not seem particularly severe.
@@
Finally, the quantile-quantile-plots based on the independent skewed $t$-distributions with mean zero, variance one, and degrees-of-freedom as well as skewness parameter values given by their estimates are presented in Figure E.3 for each residual series. They show that the marginal distributions of the series are reasonably well captured, although there is a single large outlier in the third residual series (which disappears if the March 2020 residual is removed; not shown). Therefore, the overall adequacy of the model seems, in our view, reasonable.
@@
Figure E.1: Auto- and crosscorrelation functions of the residuals of the fitted two-regime sixth-order LSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@
Figure E.2: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime sixth-order LSTVAR model for the lags $0,1,...,24$. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds ±$\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@
Figure E.3: Standardized residual time series and quantile-quantile-plots of the fitted two-regime sixth-order LSTVAR model. The quantile-quantile plots are based on $t$–distributions with zero mean, variance one, and degrees-of-freedom as well as skewness parameter values given by their estimates.

### v5

@@ Appendix E A Monte Carlo study
This appendix present a Monte Carlo study that assesses the performance of the proposed penalized maximum likelihood estimator and the three-step estimation procedure discussed in Section 4. We consider two specifications of a simple bivariate (structural) LSTVAR model with $p=1$ and $M=2$, with the first lag of the first variable as the switching variable. To maintain some constancy, we use the same impact matrices across both specifications, $\text{vec}(B_{1})=(0.6,-0.3,0.2,0.4)$ and $\text{vec}(B_{2})=(0.7,0.1,0.3,0.8)$, the same location and scale parameter values $c=0.8$ and $\gamma=5.0$, as well as the same degrees-of-freedom and skewness parameter values, $\nu_{1}=2.5$, $\nu_{2}=12.0$, $\lambda_{1}=-0.5$, and $\lambda_{2}=0.2$.
@@
We consider two specifications of the AR parameters, presented in Table E.1. In the first specification (LSTVAR 1), the AR matrices are inside the stability region (see Section 4.1). In the second specification (LSTVAR 2), the AR matrices are in the region where penalization term is strictly positive, allowing us to assess the performance of our estimation method when the true parameter value lies in the penalization region. To illustrate the closeness to the boundary, the modulus of the roots of the companion form AR matrix are presented in Table E.1. As in our empirical application, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$ for the penalization term, so the penalization term is strictly positive when at least one of the roots of the companion form AR matrix is larger than $0.95$ in modulus. The intercept parameter values are adjusted so that the unconditional means of the regimes are approximately equal in both of the specifications (see Table E.1), so that the main difference between the specifications is that the second one includes true parameter values from inside the penalization region.
@@
$\phi_{1}$ $A_{1,1}$ $\mu_{1}$ $\boldsymbol{A}_{1}||$ $\phi_{2}$ $A_{2,1}$ $\mu_{2}$ $\boldsymbol{A}_{2}||$
LSTVAR 1 $0.30$ $0.70$ 0.30-0.30 $0.00$ $0.58$ $\phantom{-}1.20$ $0.50$ $0.20$ $\phantom{-}2.00$ $0.74$
$0.60$ $0.20$ $\phantom{-}0.40$ $1.00$ $0.58$ 1.10-1.10 $0.30$ $0.50$ 1.00-1.00 $0.26$
LSTVAR 2 $0.30$ $1.10$ 0.30-0.30 $0.00$ $0.97$ $\phantom{-}0.72$ $0.74$ $0.20$ $\phantom{-}2.00$ $0.98$
$0.20$ $0.20$ $\phantom{-}0.80$ $1.00$ $0.97$ 0.87-0.87 $0.30$ $0.73$ 1.00-1.00 $0.49$
@@
Table E.1: The parameters values of $\phi_{1},\phi_{2},A_{1,1}$, and $A_{2,1}$ used for the LSTVAR models LSTVAR 1 and LSTVAR 2 in the Monte Carlo study. Also, the corresponding unconditional means of the regimes $\mu_{m}=(I_{d}-A_{m,1})^{-1}\phi_{m}$, $m=1,2$, and modulus of the eigenvalues of the companion form AR matrices of the regimes $\boldsymbol{A}_{m}||$, $m=1,2$.
@@
For each of the two model specifications, we generate $500$ samples of length $250,500,1000,2000$, and $10000$. Then, we fit the first-order two-regime LSTVAR model to each of the generated sample using the three-step estimation procedure based on the penalized log-likelihood function, proposed in Section 4. After estimation, the ordering and signs of the columns of $B_{1}$ and $B_{2}$ are standardized by assuming, without loss of generality, that the degrees-of-freedom parameters are in an ascending order and that the first skewness parameter is negative and the second one positive, in line with the true parameter values. To estimate the bias of the PML estimator, we calculate the average estimation errors $\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}$ over the $500$ Monte Carlo repetitions, and to estimate the spread of the estimates, we calculate their sample standard deviations.
@@
Table E.2: Results for the structural LSTVAR models from the Monte Carlo study assessing the performance of our penalized log-likelihood based three-step estimation method based on the PML estimator discussed in Section 4. The average estimation error is reported for the samples of length $250$, $500$, $1000$, $2000$, and $10000$ in each column first, and next to the biases are the standard deviations of the estimates in parentheses. The results are based on $500$ Monte Carlo repetitions.
@@
The results of the Monte Carlo study are presented in Table E.2. They show that there is some but mostly small bias, which diminishes when the sample size increases. Since also the standard deviation of the estimates diminishes with the sample size, our results are in line with (possible) consistency of the PML estimator. Nonetheless, for certain parameters, there appears to be a notable bias in small samples.
@@
First, the estimator of the second degrees-of-freedom parameter (true parameter value $\nu_{2}=12$) is substantially biased upwards in samples of length $T=250,500$, and $1000$, and also the standard deviation of the estimates is very large. This can explained by the property of the skewed $t$-distribution that when the degrees-of-freedom parameter value is already large, increasing it more has virtually no effect on the shape of the distribution. Consequently, if a large degrees-of-freedom parameter fits the data, the numerical optimization algorithms are incapable of appropriately discriminating between large and very large parameter values. Hence, possibly very large degrees-of-freedom parameter values drawn by random in Step 2 estimation may remain large through Step 3 estimation. However, the large upward bias does not seem concerning, as there is virtually no difference in the economic interpretations of large and very large degrees-of-freedom parameter values (with the latter amplifying the bias).
@@
Second, there is a substantial upward bias also for the scale parameter $\gamma$ when the sample size is small. A large scale parameter value indicates almost discrete regime-switches, suggesting that the PML estimator is, in small samples, biased towards estimating too fast regime-switches when they are relatively smooth in the true data generating process ($\gamma=5$). The upward bias is possibly heightened by the property that when $\gamma$ is large, the regime-switches are already nearly discrete, making it difficult for numerical estimation algorithms to properly distinguish between large and very large values, particularly in small samples.
@@
Third, the estimates of the first element of the first row of $B_{1}$ and $B_{2}$ exhibit a notable upward bias when the sample is small. A possible explanation is that the bias is related to the heavy tails of the skewed $t$-distribution, as the true degrees-of-freedom parameter value is small ($\nu_{1}=2.5$ for the related shock). In particular, the estimates may be sensitive to extreme values drawn for the heavy-tailed skewed $t$ shocks, potentially amplifying small-sample bias.
@@
In general, the small-sample bias of the PML estimator seems to be slightly larger the specification where the true AR matrices are in the penalization region (LSTVAR 2) than the specification where the true AR matrices lie well inside the stability region (LSTVAR 1). However, the differences are quite small, and even in small samples, there does not appear to be notably issues with the estimation accuracy of the AR matrices lying inside the penalization region.
Appendix F Details on the empirical application
@@
The estimation of the models, residual diagnostics, and computation of the GIRFs are carried out with the accompanying R package sstvars (Virolainen, 2025a), which implements the introduced methods. The R package sstvars also contains the dataset studied in the empirical application. The upper bound for the joint spectral radius (see Appendix B), in turn, is computed with the JSR toolbox (Jungers, 2023) in MATLAB.
@@
To study whether our model adequately captures the autocorrelation structure of the data, we depict the sample autocorrelation functions (ACF) and cross-correlation functions (CCF) of the (non-standardized) residuals for the first $24$ lags in Figure F.1. There is not much autocorrelation left in the residuals, but there are large correlation coefficients (CC) in the lag zero CCFs between the residuals of CPUI and EPUI. In addition, there are a few moderately sized CCs in the ACFs and CCFs of some of the residual series, but the vast majority of the CCs are small. Hence, our LSTVAR model appears to capture the autocorrelation structure of the data reasonably well.
@@
To study our model’s adequacy to capture the conditional heteroskedasticity in the data, we depict the sample ACFs and CCFs of squared standardized residuals for the first $24$ lags in Figure F.2. There are particularly large CC in the lag one CCFs between the IPI and EPUI, in the lag one CCF between the RATE and EPUI, and the lag zero CCF between RATE and EPUI. In addition, there are several sizeable CCs in the ACFs and CCFs of some of the residual series. Thus, there are clearly some inadequacies in capturing the conditional heteroskedasticity, but the vast majority of the CCs are small.
@@
Finally, the quantile-quantile-plots based on the independent skewed $t$-distributions with mean zero, variance one, and degrees-of-freedom and skewness parameter values given by their estimates are presented in Figure F.3 for each residual series. They show that the marginal distributions of the series are reasonably well captured, although there is a single very large outlier in the IPI’s residuals (from March 2020), which is related to the vast and fast drop in production caused by the COVID-19 lockdown. Therefore, the overall adequacy of the model seems, in our view, reasonable.
@@
Figure F.1: Auto- and crosscorrelation functions of the residuals of the fitted two-regime second-order LSTVAR model for the lags 0,1,,240,1,...,24. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds $\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@
Figure F.2: Auto- and crosscorrelation functions of the squared standardized residuals of the fitted two-regime second-order LSTVAR model for the lags 0,1,,240,1,...,24. The lag zero autocorrelation coefficients are omitted, as they are one by convention. The dashed lines are the $95\%$ bounds $\pm 1.96/\sqrt{T}$ for autocorrelations of IID observations.
@@
Figure F.3: Standardized residual time series and quantile-quantile-plots of the fitted two-regime second-order LSTVAR model. The quantile-quantile plots are based on tt–distributions with zero mean, variance one, and degrees-of-freedom and skewness parameter values given by their estimates.